Make your own problem.One problem calculating the period and the frequency of simple pendulum and one in physical pendulum.show your comple solutions to the problem.


1
Expert's answer
2021-02-10T18:23:09-0500

1)

"f=\frac{1}{2\pi}\sqrt{\frac{g}{l}}\\f=\frac{1}{2\pi}\sqrt{\frac{9.8}{1}}=0.498\ Hz\\T=2\pi\sqrt{\frac{1}{9.8}}=2.01\ s"

2)


"f=\frac{1}{2\pi}\sqrt{\frac{mgh}{I}}\\f=\frac{1}{2\pi}\sqrt{\frac{(1)(1)9.8}{1}}=0.498\ Hz\\T=2\pi\sqrt{\frac{1}{(1)(1)9.8}}=2.01\ s"